OpenStudy (anonymous):

X^-2 = (2+or-square root of 3) Drawn out please

OpenStudy (anonymous):

I showed you already. Please be patient and don't repost. Did you have questions about the solution I presented?

OpenStudy (anonymous):

i need to see it worked out to get it

OpenStudy (anonymous):

im new at this but is there a way to draw it out to see it?

OpenStudy (anonymous):

\[x^{-2}=2 \pm \sqrt3\]\[\frac{1}{x^2}=2 \pm \sqrt3\]\[x^2=\frac1{2 \pm \sqrt3}\]\[x=\sqrt{\frac1{2 \pm \sqrt3}}\]

OpenStudy (anonymous):

Just look at my solution one step at a time

OpenStudy (anonymous):

If you are confused about a step, ask what was done or why

OpenStudy (anonymous):

each line is a single step

OpenStudy (anonymous):

\[\large x^{-2} = 2 \pm \sqrt{3}\]\[\large \implies \frac{1}{x^2} = 2\pm \sqrt{3}\]\[\large \implies x^2 = \frac{1}{2\pm \sqrt{3}}\]\[\large \implies x^2 = \frac{1}{2\pm \sqrt{3}} \cdot \frac{2 \mp \sqrt{3}}{2\mp \sqrt{3}}\]\[\large \implies x^2 = \frac{2\pm \sqrt{3}}{4-3} \]\[\large \implies x^2 = \frac{2 \pm \sqrt{3}}{1} \]\[\large \implies x^2 = 2 \pm \sqrt{3}\]\[\large \implies x = \pm \sqrt{2 \pm \sqrt{3}}\]

OpenStudy (anonymous):

how did you go from ⟹x2=2±√3/4−3 ⟹x2=2±√3/1

OpenStudy (anonymous):

Which line?

OpenStudy (anonymous):

how did you go from 5 to 6

OpenStudy (anonymous):

4-3 = 1

OpenStudy (anonymous):

All I did was simplify the denominator.

OpenStudy (anonymous):

oh wow im dumb idk how i missed that

OpenStudy (anonymous):

I think the more salient question is the step from 4 to 5

OpenStudy (anonymous):

Did you understand what happened there?

OpenStudy (anonymous):

yeah i understood that